# A Multi-Component, Multi-Physics Computational Model for Solving Coupled Cardiac Electromechanics and Vascular Haemodynamics

### v1

@@ 2.3 Coupled Model
consisting of the gathering of equations 13 and 48. Similarly, we compactly denote the vascular hemodynamics model — solved in HemeLB — as
@@
Analogously, velocity continuity is enforced by assuming that the decreased rate of ventricular volume $V_{\mathrm{LV}}$ equals the inlet flow rate $Q_{\mathrm{in}}$ across $\Gamma_{\mathrm{in}}^{\mathcal{H}}$. This kinematic continuity is formulated as
$Q_{\mathrm{in}}(t)=-\frac{\mathrm{d}V_{\mathrm{LV}}(\bm{d})}{\mathrm{d}t}\quad\text{in }\Gamma_{\mathrm{in}}^{\mathcal{H}}\times[t_{0},t_{f}].$ (22d)
@@ 2.4 Partitioned Coupling Scheme between the Two Models
Alya and HemeLB are coupled via a partitioned scheme, wherein their respective model equations are solved independently using exchanged data. During intervals between data exchanges, Alya and HemeLB independently solve their model equations 22a and 22b using the exchanged data $p_{\mathrm{endo}}$ and $Q_{\mathrm{in}}$ as the boundary values on the coupling interfaces, respectively. In addition, Alya computes the flow rate $Q_{\mathrm{in}}$ on $\Gamma^{\mathcal{A}}_{0_{\mathrm{endo}}}$ and transmits it to HemeLB, while HemeLB determines the pressure $p_{\mathrm{endo}}$ on $\Gamma_{\mathrm{in}}^{\mathcal{H}}$ and shares it with Alya.
@@
Since Alya and HemeLB focus on different dynamic scales, the time step sizes required for the simulations differ. Therefore, allowing the coupling to occur at various frequencies in the two solvers is important. This capability is implemented in our algorithm by controlling the parameter $k$, defined as the ratio of the time step size of Alya to that of HemeLB, i.e. $k=\Delta t^{\mathcal{A}}/\Delta t^{\mathcal{H}}$. This ratio is greater than or equal to one because the electromechanics model in Alya operates on a larger dynamic scale than the mesoscopic fluid dynamics scale in HemeLB. In addition, this ratio is defined as an integer so that the boundary value in Alya is updated at every time step. Let $m\in[0,k-1]$ be an iteration index. Each time HemeLB completes a time step, $m$ is incremented by one. When $m$ reaches $k$, HemeLB updates the value of $Q_{\mathrm{in}}$ from File B and resets $m$ to zero. This ensures that data exchange occurs once per time step of Alya and once every $k$ time steps of HemeLB. The algorithm for this coupling scheme is outlined in Algorithm 1.
@@ 2.5 Simulation Tests
We conduct three simulation tests to study different aspects of the proposed coupled model. Test 0 sets the baseline solution of the heart electromechanics for subsequent comparisons with the coupled model. Test 1 verifies the coupling scheme using simplified geometries. Test 2 demonstrates the feasibility of the coupled model within physiological contexts.
2.5.1 Test 0: Left Ventricle
In this test, we perform a standalone simulation using Alya. We use the results of this simulation to (i) determine the reference pressure $P_{\text{ref}}$ of the fluid domain of HemeLB and (ii) compare the results with those obtained from the coupled Alya-HemeLB simulations. The geometry for this simulation represents the left ventricle, which is based on a 3D model of the human heart produced from computed tomography and magnetic resonance imaging scans and created by Zygote [76]. The endocardial pressure during the ejection phase is determined by the WK2 model, as detailed in Section 2.1. The time step size $\Delta t^{\mathcal{A}}$ is set to $10^{-4}$ s for both electrophysiology and mechanics. The parameters of the physics models used in Alya are given in Table 1.
@@ 2.5.1 Test 0: Left Ventricle
$I_{\mathrm{app}}$ Applied current $\mu$A/cm2 -40
$C_{\mathrm{m}}$ Membrane capacitance per unit area $\mu$F/cm2 1
@@
$C^{\mathcal{A}}$ Compliance of Windkessel model for ejection ml/barye 0.001
$R^{\mathcal{A}}$ Resistance of Windkessel model for ejection barye s / ml 1500
@@
2.5.2 Test 1: Left Ventricle with Cylinder
In the first test, we verify the coupling scheme, focusing on its implementation and computational aspects. Specifically, we check whether the coupling occurs at the correct time steps and whether the values exchanged at the coupling interface are consistent in Alya and HemeLB. The simulation domains include the same left ventricle geometry as in Test 0 and a cylinder representing the aorta.
A physiological cardiac output poses tight constraints on the parameters of the fluid flow simulation. First, for a given flow rate, the flow speed is inversely proportional to the area of the inlet. Second, according to equation 12, the Mach number is proportional to the flow speed, the time step size, and the inverse of the voxel size. However, the LBGK model requires that the non-dimensionalised relaxation time, $\tau_{0}/\Delta t^{\mathcal{H}}$, is not close to 0.5 for numerical stability [60]. According to equation 11, the dynamic viscosity is proportional to the square of the voxel size and the inverse of the time step size. These relations result in the competing conditions that
@@ 2.5.2 Test 1: Left Ventricle with Cylinder
    increasing the voxel size will decrease the Mach number and the relaxation time;
@@
    increasing the time step size will increase the Mach number and the relaxation time; and
@@
    increasing the inlet area will increase the Mach number but decrease the relaxation time.
@@
The above constraints lead to the following choice of parameters. A cylinder of radius 0.02 m and a height of 0.2 m is used, as shown in Figure 5. It is voxelised to a resolution ($\Delta x^{\mathcal{H}}$) of $1.67\cdot 10^{-4}$ m so that there are about 120 lattice sites along a radius. The time step size $\Delta t^{\mathcal{H}}$ is varied between $2.5\cdot 10^{-5}$ s and $1.0\cdot 10^{-5}$ s to check the implementation of the coupling scheme. These values correspond to a coupling frequency of $k=\{4,5,8,10\}$ given the time step size of $10^{-4}$ s for Alya. In particular, we study the solutions obtained from the simulation with $\Delta t^{\mathcal{H}}=2.5\cdot 10^{-5}$ s, or $\tau_{0}/\Delta t^{\mathcal{H}}=0.5108$. For a physiological cardiac output resulting in a maximum flow rate of around 0.5 dm3/s, the maximum velocity assuming a quadratic profile is 0.796 m/s. The corresponding Mach number of 0.208 is considered to contribute to a small compressibility error [60].
In addition, the resistance and capacitance parameters of the WK2 model at the outlet are set to $1.5\cdot 10^{8}$ Pa s/m3 and $10^{6}$ m3/Pa, respectively. The viscous sponge layer enclosing the outlet is set to have a radius of $R_{v}=0.04$ m, which is two times that of the cylinder.
@@
2.5.3 Test 2: Left Ventricle with Thoracic Aorta
In the second test, we evaluate the coupled model in a physiologically sound setting. We use the same left ventricle model for Alya as in the first test but employ a thoracic aorta model for HemeLB. This arterial model, shown in Figure 3, is adapted from the 0012_H_AO_H model in the Vascular Model Repository [66], which was produced from a computed tomography angiogram of a 26-year-old male human. The model includes the ascending aorta, the descending aorta, and three branches: brachiocephalic, left common carotid, and left subclavian arteries. The original model is scaled up by 1.25 times in all dimensions to ensure the inlet radius approximates the normal size for adults (0.015 m [77, 78]). In addition, the inlet and outlets are elongated by one to two times their diameters to mitigate instabilities caused by the boundary conditions [79].
Table 2: Parameters of the two-element Windkessel model imposed at the outlets of the thoracic aorta model.
Outlet 0 Outlet 1 Outlet 2 Outlet 3
Resistance (kg m-4 s-1) $3.422\cdot 10^{9}$ $4.182\cdot 10^{9}$ $1.176\cdot 10^{9}$ $2.543\cdot 10^{8}$
Capacitance (m4 s2 kg-1) $3.443\cdot 10^{-10}$ $2.817\cdot 10^{-10}$ $1.002\cdot 10^{-9}$ $4.632\cdot 10^{-9}$
The fluid domain for HemeLB is created by voxelising [80] the thoracic aorta model to obtain a three-dimensional uniform grid of lattice sites. A voxel size of $\Delta x^{\mathcal{H}}=10^{-4}$ m is used, resulting in about 150 lattice sites along a radius of the inlet. The time step size $\Delta t^{\mathcal{H}}$ is set to $6.25\cdot 10^{-6}$ s, or $\tau_{0}/\Delta t^{\mathcal{H}}=0.50625$, which corresponds to a coupling frequency of $k=16$ given the time step size of $10^{-4}$ s for Alya. For a physiological cardiac output resulting in a maximum flow rate of around 0.5 dm3/s, the maximum velocity assuming a quadratic profile is 1.42 m/s. The corresponding Mach number of 0.153 is considered to contribute to a small compressibility error [60].
The low-pass filter with an appropriate smoothing factor $\alpha$ as described in equation 24 is applied to stabilise the simulation. The parameters of the WK2 model at the outlets are obtained using the method we introduced previously [70, 81] and the flow rate ratios $5.5:4.5:16:74$ for Outlets 0 to 3, respectively [82]. The resulting parameters, listed in Table 2, have the same order of magnitude as those found in similar studies of the aorta [83, 84]. The radius of the viscous sponge layer, $R_{v}$, is set to 0.03 m at all outlets so that it encloses two times the largest outlet, which has an equivalent radius of 0.0115 m.

### v2

@@ 2.3 Coupled Model
consisting of the gathering of equations 13 and 49. Similarly, we compactly denote the vascular hemodynamics model — solved in HemeLB — as
@@
Analogously, velocity continuity is enforced by assuming that the decreased rate of ventricular volume $V$ equals the inlet flow rate $Q_{\mathrm{in}}$ across $\Gamma_{\mathrm{in}}^{\mathcal{H}}$. This kinematic continuity is formulated as
$Q_{\mathrm{in}}(t)=-\frac{\mathrm{d}{\color[rgb]{0,0,0}V}(\bm{d})}{\mathrm{d}t}\quad\text{in }\Gamma_{\mathrm{in}}^{\mathcal{H}}\times[t_{0},t_{f}].$ (25d)
The flow rate is computed using backward finite differences with the volume in equation 10.
@@ 2.4 Partitioned Coupling Scheme between the Two Models
Alya and HemeLB are coupled via a partitioned scheme, wherein their respective model equations are solved independently using exchanged data. During intervals between data exchanges, Alya and HemeLB independently solve their model equations 25a and 25b using the exchanged data $p_{\mathrm{endo}}$ and $Q_{\mathrm{in}}$ as the boundary values on the coupling interfaces, respectively. In addition, Alya computes the flow rate $Q_{\mathrm{in}}$ on $\Gamma^{\mathcal{A}}_{0_{\mathrm{endo}}}$ and transmits it to HemeLB, while HemeLB determines the pressure $p_{\mathrm{endo}}$ on $\Gamma_{\mathrm{in}}^{\mathcal{H}}$ and shares it with Alya.
@@
Since Alya and HemeLB focus on different dynamic scales, they require different time step sizes. To accommodate this difference, the coupling scheme allows data exchange at different frequencies in the two solvers, controlled by the parameter $k$. This parameter is defined as the ratio of the time step size of Alya to that of HemeLB, i.e. $k=\Delta t^{\mathcal{A}}/\Delta t^{\mathcal{H}}$, and must be an integer greater than or equal to one. This constraint ensures that Alya, which models electromechanics on a larger dynamic scale, updates its boundary conditions at every time step, while HemeLB, operating at a mesoscopic fluid dynamics scale, updates less frequently. To manage data exchange, an iteration index $m\in[0,k-1]$ is introduced. Each time HemeLB completes a time step, $m$ is incremented by one. When $m$ reaches $k$, HemeLB retrieves the updated value of $Q_{\mathrm{in}}$ from File B and resets $m$ to zero. This ensures that data exchange occurs once per time step of Alya and once every $k$ time steps of HemeLB.
The algorithm for this coupling scheme is outlined in Algorithm 1.
@@
While implicit coupling with sub-iteration is a viable option to improve the stability of the coupled model, it comes with a significant increase in computational cost. Given the novelty of our 3D multi-component, multi-physics cardiovascular model, our focus in this study is to establish its feasibility and demonstrate its potential for medical applications before exploring additional coupling techniques. In this study, we employ an explicit coupling scheme for its computational efficiency and simplicity.
@@ 2.5 Simulation Tests
We conduct four simulation tests to study different aspects of the proposed coupled model. Test 0 sets the baseline solution for both cardiac electromechanics and vascular haemodynamics for subsequent comparisons with the coupled model. Test 1 verifies the coupling scheme using simplified geometries. Test 2 demonstrates the feasibility of the coupled model within physiological contexts and highlights its differences from the standalone simulations in Test 0. Test 3 applies the coupled model to study how myocardial scarring affects downstream vascular flow.
2.5.1 Test 0: Baseline Solutions for Left Ventricle and Thoracic Aorta
In this test, we perform standalone simulations using Alya and HemeLB separately. These simulations establish the baseline solutions for comparisons with the coupled Alya-HemeLB simulations in Test 2. In addition, the results of the Alya simulation determine the reference pressure $P_{\text{ref}}$ in the fluid domain of HemeLB (see equation 18).
@@ 2.5.1 Test 0: Baseline Solutions for Left Ventricle and Thoracic Aorta
$I_{\mathrm{app}}$ Applied current µA/cm2 -40
$C_{\mathrm{m}}$ Membrane capacitance per unit area µF/cm2 1
@@
$C^{\mathcal{A}}$ Compliance of Windkessel model for ejection ml/barye 0.0005
$R^{\mathcal{A}}$ Resistance of Windkessel model for ejection barye s / ml 2700
@@
The geometry for the Alya simulation represents the left ventricle, which is based on a 3D model of the human heart produced from computed tomography and magnetic resonance imaging scans and created by Zygote [78]. The endocardial pressure during the ejection phase is determined by the WK2 model, as detailed in Section 2.1. The time step size $\Delta t^{\mathcal{A}}$ is set to $10^{-4}$ s for both electrophysiology and mechanics. The parameters of the physics models used in Alya are given in Table 1.
For the simulation performed with HemeLB, we employ a model of the thoracic aorta. This arterial model, shown in Figure 3, is adapted from the 0012_H_AO_H model in the Vascular Model Repository [66], which was produced from a computed tomography angiogram of a 26-year-old male human. The model includes the ascending aorta, the descending aorta, and three branches: brachiocephalic, left common carotid, and left subclavian arteries. The original model is scaled up by 1.25 times in all dimensions to ensure the inlet radius approximates the normal size for adults (0.015 m [79, 80]). In addition, the inlet and outlets are elongated by one to two times their diameters to mitigate instabilities caused by the boundary conditions [81].
The fluid domain for HemeLB is created by voxelising [82] the thoracic aorta model to obtain a three-dimensional uniform grid of lattice sites. A voxel size of $\Delta x^{\mathcal{H}}=10^{-4}$ m is used, resulting in about 150 lattice sites along a radius of the inlet and 253 million lattice sites in the entire simulation domain. The time step size $\Delta t^{\mathcal{H}}$ is set to $6.25\cdot 10^{-6}$ s, or $\tau_{0}/\Delta t^{\mathcal{H}}=0.50625$. This corresponds to a coupling frequency of $k=16$ given the time step size of $10^{-4}$ s for Alya when the models are coupled in Test 2. For a physiological cardiac output resulting in a maximum flow rate $Q_{\mathrm{in}}$ of around 0.45 dm3/s [83], the maximum velocity assuming a quadratic profile is 1.27 m/s. The corresponding Mach number of 0.138 is considered to contribute to a small compressibility error [60].
Table 2: Parameters of the three-element Windkessel model imposed at the outlets of the thoracic aorta model.
Outlet 0 Outlet 1 Outlet 2 Outlet 3
$R_{c}$ (kg m-4 s-1) $2.050\cdot 10^{8}$ $2.500\cdot 10^{8}$ $7.050\cdot 10^{7}$ $3.040\cdot 10^{7}$
$R_{p}$ (kg m-4 s-1) $3.422\cdot 10^{9}$ $4.182\cdot 10^{9}$ $1.176\cdot 10^{9}$ $2.543\cdot 10^{8}$
$C_{p}$ (m4 s2 kg-1) $3.443\cdot 10^{-10}$ $2.817\cdot 10^{-10}$ $1.002\cdot 10^{-9}$ $4.632\cdot 10^{-9}$
The low-pass filter with an appropriate smoothing factor $\alpha$ as described in equation 27 is applied to stabilise the simulation. To determine the parameters of the WK3 models at the outlets, we first estimate $R_{p}$ and $C_{p}$ using the method we introduced previously [70, 84] with the flow rate ratios $5.5:4.5:16:74$ for Outlets 0 to 3, respectively [85]. After that, we set $R_{c}$ by assuming $R_{c}/(R_{c}+R_{p})=5.6\%$, following the approach of Bonfanti et al. [86], and then fine-tune its value. The resulting parameters are listed in Table 2. The radius of the viscous sponge layer, $R_{v}$, is set to 0.03 m at all outlets so that it encloses two times the largest outlet, which has an equivalent radius of 0.0115 m.
2.5.2 Test 1: Coupled Simulations of Left Ventricle and Cylinder
In the first test, we verify the coupling scheme, focusing on its implementation and computational aspects. Specifically, we check whether the coupling occurs at the correct time steps and whether the values exchanged at the coupling interface are consistent in Alya and HemeLB and perform a convergence study. The simulation domains include the same left ventricle geometry as in Test 0 and a cylinder representing the aorta.
A physiological cardiac output poses tight constraints on the parameters of the fluid flow simulation. First, for a given flow rate $Q_{\mathrm{in}}$, the flow velocity $\bm{U}$ is inversely proportional to the inlet area. Second, from equation 14, the Mach number is proportional to the flow velocity $\bm{U}$, the time step size $\Delta t^{\mathcal{H}}$, and the inverse of the voxel size $\Delta x^{\mathcal{H}}$. Third, from equation 13, for a given kinematic viscosity $\nu_{0}$, the non-dimensional relaxation time $\tau_{0}/\Delta t^{\mathcal{H}}$ scales linearly with the time step size and inversely with the square of the voxel size:
$\frac{\tau_{0}}{\Delta t^{\mathcal{H}}}=3\nu_{0}\frac{\Delta t^{\mathcal{H}}}{(\Delta x^{\mathcal{H}})^{2}}+\frac{1}{2}.$ (28)
These relations result in the competing conditions that
@@ 2.5.2 Test 1: Coupled Simulations of Left Ventricle and Cylinder
    decreasing the inlet area will increase the Mach number;
@@
    decreasing the voxel size will increase the Mach number and the relaxation time; and
@@
    decreasing the time step size will decrease the Mach number and the relaxation time.
However, the Mach number must remain below 0.3 to minimise the compressibility errors [60]. Additionally, for numerical stability, the LBGK model requires that $\tau_{0}/\Delta t^{\mathcal{H}}$ does not approach 0.5 [60].
@@
Considering the above constraints, we adopt the following choice of parameters. A cylinder of radius 0.02 m and a height of 0.4 m is used, as shown in Figure 5. It is voxelised to a resolution ($\Delta x^{\mathcal{H}}$) of $1.67\cdot 10^{-4}$ m, resulting in about 120 lattice sites along a radius and 108 million lattice sites in the entire simulation domain. The time step size $\Delta t^{\mathcal{H}}$ is varied between $2.5\cdot 10^{-5}$ s and $1.0\cdot 10^{-5}$ s to check the implementation of the coupling scheme. These values correspond to coupling frequencies of $k=\{4,5,8,10\}$ given the time step size of $10^{-4}$ s for Alya. In particular, we study the solutions obtained from the simulation with $\Delta t^{\mathcal{H}}=2.5\cdot 10^{-5}$ s, which corresponds to $\tau_{0}/\Delta t^{\mathcal{H}}=0.509$. For a physiological cardiac output resulting in a maximum flow rate $Q_{\mathrm{in}}$ of around 0.45 dm3/s [83], the maximum velocity assuming a quadratic profile is 0.716 m/s. The corresponding Mach number of 0.186 is considered to contribute to a small compressibility error [60].
In addition, the parameters of the WK3 model at the outlet are set to $R_{c}=2.0\cdot 10^{6}$ kg m-4 s-1, $R_{p}=1.0\cdot 10^{9}$ kg m-4 s-1, and $1.5\cdot 10^{-8}$ m4 s2 kg-1. The viscous sponge layer enclosing the outlet is set to have a radius of $R_{v}=0.04$ m, which is two times that of the cylinder.
@@
Moreover, we perform a convergence study to assess the impact of grid resolution on the solutions of the coupled model. To evaluate error convergence, we compare the pressure and flow rate at the inlet at the ejection peak (0.865 s), using the solutions on the finest grid as a reference. The absolute differences from these reference solutions are denoted as $\varepsilon_{P}$ for the pressure and $\varepsilon_{Q}$ for the flow rate.
Grid refinement for lattice Boltzmann simulations requires a coordinated reduction of the time step size ($\Delta t^{\mathcal{H}}$) and lattice spacing ($\Delta x^{\mathcal{H}}$) to ensure numerical stability and physical consistency. Two principal scaling approaches exist [60]: acoustic scaling maintains a constant $\Delta t^{\mathcal{H}}/\Delta x^{\mathcal{H}}$ ratio to preserve the Mach number (see equation 14) and wave propagation characteristics, while diffusive scaling maintains a constant $\Delta t^{\mathcal{H}}/(\Delta x^{\mathcal{H}})^{2}$ ratio to preserve the non-dimensional relaxation time $\tau_{0}/\Delta t^{\mathcal{H}}$ and kinematic viscosity $\nu_{0}\Delta t^{\mathcal{H}}/(\Delta x^{\mathcal{H}})^{2}$ (see equation 13). We adopt acoustic scaling because it ensures a consistent representation of pulsatile flow phenomena across refinement levels. In addition, we maintain the coupling frequency $k=4$, which is not a grid parameter, by proportionally reducing $\Delta t^{\mathcal{A}}$ while keeping $\Delta x^{\mathcal{A}}$ fixed. The convergence properties of Alya under grid refinement have been analysed in previous standalone studies [87, 5], demonstrating its numerical accuracy across a range of spatial resolutions.
Due to the substantial memory demands of voxelisation, generating the cylinder domain with an arbitrarily fine spatial resolution is impractical. To obtain more data points while managing computational constraints, we adopt a refinement ratio of 0.75 instead of the commonly used 0.5 (grid doubling). This choice remains viable, as we compare spatially averaged quantities (i.e. time-dependent only) rather than pointwise values across different resolutions.
2.5.3 Test 2: A Coupled Simulation of Left Ventricle and Thoracic Aorta
In the second test, we evaluate the coupled model in a physiologically sound setting and assess the impact of coupling on the individual models. We use the same simulation setup as in Test 0 with the left ventricle geometry for Alya and the thoracic aorta geometry for HemeLB and couple the two codes. The coupling frequency is set to $k=16$ with $\Delta t^{\mathcal{A}}=10^{-4}$ s and $\Delta t^{\mathcal{H}}=6.25\cdot 10^{-6}$ s.
To assess the impact of coupling on the simulated outcomes, we compare the displacement of heart muscles, $\bm{d}$, and the time-averaged wall shear stress (TAWSS) in the aorta between the coupled simulation and the standalone Alya and HemeLB simulations in Test 0. We compute the percentage differences in these quantities using the following formula:
[…]
TAWSS is a key biomechanical parameter that quantifies the shear forces exerted by blood flow on vessel walls. Abnormal TAWSS levels are associated with the progression of various vascular diseases, including aneurysms, atherosclerosis, and thrombosis [88, 89]. Accurately capturing TAWSS is therefore essential for characterising the haemodynamic environment and evaluating potential risks. The computation of TAWSS in HemeLB is described in our previous work [84].
2.5.4 Test 3: Impact of Myocardial Scarring on Downstream Vascular Flow
Figure 6: Left ventricle with idealised scars of different sizes: $r_{\mathrm{scar}}=$ 0.5 cm, 1.0 cm, and 2.5 cm (left, middle, and right panels, respectively).
As a final test, we evaluate whether the model can capture physiologically relevant features of disease. Specifically, we investigate the impact of myocardial scarring in the left ventricle on aortic haemodynamics by repeating the coupled simulation from Test 2 under pathological conditions. To this end, we generate idealised scars of varying sizes, as illustrated in Figure 6. Scars are created by selecting a point on the endocardial surface, computing the Euclidean distance from this point to each mesh node, and applying a threshold ${r}_{\mathrm{scar}}$ of 0.5 cm, 1.0 cm, and 2.5 cm to define the scar region. The scarred regions were considered electrically non-conductive: the cellular model is deactivated ($I_{\mathrm{ion}}=0$ in equation 1), and the monodomain diffusivity $\bm{D}$ is set to zero [90, 91]. By varying the size of the scar, we aim to assess how different degrees of cardiac dysfunction influence haemodynamic outcomes. In particular, we compare the TAWSS distributions in the aorta with the distribution of Test 2 using the percentage difference defined in equation 29.